# Optimizing Multi-Modal Trackers via Sensitivity-aware Regularized Tuning

### v1

@@ Abstract
This paper tackles the critical challenge of optimizing multi-modal trackers by effectively adapting the pre-trained models for RGB data. Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off. To mitigate this dilemma, we propose a novel sensitivity-aware regularized tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities. Through a comprehensive investigation from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are primary drivers of this issue. Specifically, we first analyze the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization. Then, we further explore transfer sensitivities during the tuning phase, emphasizing adaptability and stability. By incorporating these sensitivities as regularization terms, our method significantly enhances the transferability across modalities. Extensive experiments showcase the superior performance of the proposed method, surpassing current state-of-the-art techniques across various multi-modal tracking. The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
Index Terms:
Multi-modal Tracking, Cross-modal Transfer, Parameter Sensitivity, Regularized Tuning.
Figure 1: Training and testing performance on LasHeR, comparing our method with existing fine-tuning approaches. Our method effectively mitigates the misfitting issue and enhances both generalization and stability of the multi-modal tracker.
@@ I Introduction
Object tracking, a foundation task of visual perception, has seen significant advancements over the past decades [1, 2, 3, 4]. Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions. Therefore, multi-modal tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest. With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [5, 6, 7, 8]. There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
Some existing approaches have explored the full fine-tuning (FFT) paradigm [9, 10, 11], where these trackers are initialized with pre-trained weights and subsequently tuned for task-specific objectives. Nevertheless, due to the significant gap and limited scale of auxiliary modalities, they are intractable to retain the pre-trained knowledge during transfer and tend to induce catastrophic forgetting and overfitting (i.e., widened train-test gap). In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [12, 13, 14]. The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modal-specific ones to retain pre-trained knowledge. Several methods fall under this umbrella [15, 16, 17], including prompt tuning, visual adapter, etc. While effective, PEFT methods impose excessive constraints on the primary model weights, resulting in underfitting (i.e., restricted train upper bound) for handling the vast distribution drift. In summary, existing tuning paradigms fluctuate between excessive flexibility and overly rigid constraints, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and adaptation (please refer to Fig. 1).
In this work, we endeavor to mitigate the misfitting issue in cross-modal tracker adaptation by strategically regularizing the learning process. By examining parameter responses spanning from the pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability. To mitigate this, we propose a sensitivity-aware regularized tuning framework that delicately calibrates the gradient updates to boost model transfer with precision. Specifically, we optimize multi-modal trackers from the following perspectives.
@@
    Formulating Prior Sensitivity. We begin by investigating the tangent space of pre-trained parameters, using it as a critical indicator to assess and preserve prior generalizability. Moreover, we estimate such prior sensitivity via an eigen-decomposition approximation.
@@
    Modeling Transfer Sensitivity. To further elucidate transfer challenges, we explore how sparse gradients exacerbate tuning instability. Accordingly, we construct this transfer sensitivity using the off-the-shelf gradient matrix, with the aim of facilitating gradient rebalancing.
@@
    Sensitivity-Aware Regularization Tuning. By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modal representations through finely modulated, sensitivity-driven updates. This regularization tuning facilitates seamless transfer to various multi-modal tracking tasks, which can continuously enhance the model in the training phase.
Our method strategically guides the cross-modal fine-tuning process to optimize downstream tracking tasks. Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modal tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models. Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
@@
  • we revisit the misfitting issue of multi-modal tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of multi-modal trackers;
  • we conduct comprehensive experiments covering three multi-modal tracking tasks and seven datasets using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels;
The remainder of the paper is organized as follows. Section II reviews existing works on multi-modal trackers and cross-modal transfer learning methods. In Section III, we explicitly formulate the prior and transfer sensitivity from both the pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-aware regularized tuning. In Section IV, we present experiments on seven multi-modal tracking benchmarks to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs. Finally, Section V concludes the paper.
@@ II Related Work
II-A Multi-Modal Tracking
Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [18, 19, 20, 21]. In recent years, due to the unreliability of RGB data in challenging scenarios, increasing studies are expected to integrate auxiliary modalities (e.g., event, depth, thermal data) to enhance tracking performance. These complementary signals compensate for the inherent limitations of RGB imagery [22]. For instance, event cameras provide robust dynamic scene information under extreme motion or lighting. Leveraging this, [23, 24, 25, 26, 27, 28] combine RGB and event streams to predict objects in high-speed and low-dynamic scenarios. Meanwhile, depth data offers spatial cues that are particularly useful for handling occlusion and background clutter. [29, 30, 31, 32] exploit this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes. Similarly, thermal infrared imaging captures heat signatures independent of visible light, proves effective in low-illumination. [33, 34, 35, 36] demonstrate that fusing thermal and RGB data can yield more reliable appearance representations.
In summary, these methods emphasize effective multi-modal interaction and fusion [37, 9, 38, 39]. With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [5, 40, 8] have achieved remarkable generalization across diverse scenarios. These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers. Multi-modal tracking is increasingly driven by reusing RGB pre-trained semantic priors, particularly when annotated multi-modal data are scarce or noisy. In this work, we target this challenge by focusing on optimizing the adaptation of pre-trained trackers for efficient cross-modal transfer.
@@ II-B Cross-modal Transfer Learning
To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged. Some works follow the full fine-tuning (FFT) paradigm [37, 9, 10, 11], where the pre-trained models serve as initializations and are entirely re-trained on the target tasks. These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model. Representatively, SUTrack [8] employs a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization. While often effective, one primary drawback may be innate to this paradigm: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation often leads to significant catastrophic forgetting or overfitting. To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [13, 14, 12], some parameter-efficient fine-tuning (PEFT) methods have been proposed. The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task. By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting. For instance, ProTrack [41] and ViPT [15] modulate RGB features by introducing trainable auxiliary tokens into attention layers. Similarly, methods like BAT [42] and SDSTrack [16] insert lightweight adapter modules between attention layers for cross-modal shift compensation. More recently, unified trackers like UnTrack [17] and OneTracker [1] employ LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs. Despite their effectiveness, excessive constraints on pre-trained weights can incur insufficient adaptation. Furthermore, the added modules tend to over-fit since the additional parameters are optimized from scratch when tailored to modality-specific tasks. Consequently, how to efficiently fine-tune RGB-based pre-trained models for multi-modal tracking remains a significant challenge.
Figure 2: Model architecture for our multi-modality tracker.
@@ III Proposed Method
Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers. To unlock the full potential of pre-trained trackers, we revisit the core principles of cross-modal transfer process. In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning. Next, in Section III-B, we identify pronounced prior and transfer sensitivities that contribute to catastrophic forgetting and training instability, and further quantify these critical sensitivities. Finally, in Section III-C, we incorporate these sensitivities into the tuning phase, enabling adaptive and dynamic modulation of parameter updates.
@@ III-A Preliminaries
Network Architecture. Fig. 2 depicts the overall architecture of our multi-modal tracker. The RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens. Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction. Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens. To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
Cross-Domain Transfer. We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameter $\theta$. It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box. In this paper, we focus on efficiently transferring the models pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$. In this setting, let $\theta_{0}$ be the pre-trained parameters (optimal on $D_{0}$), which serve as initialization. After fine-tuning on $D_{t}$, the parameters become $\theta_{t}$. Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
$\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\theta^{(0)}=\theta_{0},$ (1)
However, directly optimizing $\theta_{t}$ as per Eq. (1) leads to severe catastrophic forgetting and unstable tuning, as evidenced by the pronounced over-fitting and oscillations in training dynamics observed in Fig. 1. This motivates a more principled transfer learning approach described next.
Figure 3: Loss-parameter manifold schematic for different fine-tuning paradigms. The “FFT” fine-tunes all pre-trained weights without regularization, leading to significant catastrophic forgetting on pre-trained dataset. Although “PEFT”-based methods only tune the additional parameters to maintain the pre-traiend knowledge, but their coarse modeling limits performance on new domain. In contrast, our “SRFT” method carefully optimize the model within an approximated tangent space of the pre-trained manifold, indicating a better plasticity-stability trade-off between pre-trained knowledge and adaptation.
III-B Modeling Sensitivity for Multi-Modal Tracking

### v2

# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

@@ Abstract
This paper tackles the critical challenge of optimizing multi-modality trackers by effectively adapting pre-trained models for RGB data. Existing fine-tuning paradigms oscillate between excessive freedom and over-restriction, both leading to a suboptimal plasticity-stability trade-off. To mitigate this dilemma, we propose a novel sensitivity-regularized fine-tuning framework, which delicately refines the learning process by incorporating intrinsic parameter sensitivities. Through a comprehensive investigation of the transition from pre-trained to multi-modal contexts, we identify that parameters sensitive to pivotal foundational patterns and cross-domain shifts are the primary drivers of this issue. Specifically, we first probe the tangent space of pre-trained weights to measure and orient prior sensitivities, dedicated to preserving generalization. Subsequently, we characterize transfer sensitivities during the tuning phase, emphasizing adaptability and stability. By incorporating these sensitivities as unified regularization terms, our method significantly enhances the transferability across modalities. Extensive experiments showcase the superior performance of our method, surpassing current state-of-the-art techniques across various multi-modality tracking benchmarks. The source code and models will be publicly available at https://github.com/zhiwen-xdu/SRTrack.
Figure 1: Optimization trajectory analysis on LasHeR. This plot contrasts the training and testing dynamics of different tuning paradigms. As visualized by the colorful arrows, our method effectively mitigates the misfitting issue and enhances multi-modality trackers with superior generalization and stability.
@@ I Introduction
Object tracking, a foundation task of visual perception, has seen significant advancements over the past decades [16, 54, 67, 4]. Despite the promising results, RGB-based trackers often struggle with some complex and degraded conditions, such as extreme illumination, motion blur, and occlusions. Therefore, multi-modality tracking with more comprehensive information (e.g., event, depth, thermal) has garnered growing interest. With the popularity of the data-driven methods in the object tracking community, both data scale and model size have experienced huge explosions [58, 29, 10, 9]. There is a prevailing paradigm that explores pre-trained trackers on large-scale RGB-based datasets and adapts them to diverse auxiliary modalities, a process known as cross-modal fine-tuning or transfer learning, to enhance performance and accelerate convergence.
Some existing approaches have explored the full fine-tuning (FFT) paradigm [45, 73, 41], where trackers are initialized with pre-trained weights and subsequently optimized for task-specific objectives. Nevertheless, due to the significant domain gaps and limited scale of auxiliary modalities, these methods struggle to preserve the pre-trained knowledge structure during transfer, often inducing severe overfitting (i.e., a widened train-test gap and deteriorating test performance). In contrast to full fine-tuning, recent research has shifted toward parameter efficient fine-tuning (PEFT) [18, 21, 7]. The PEFT keeps the majority of pre-trained parameters frozen, updating only a small fraction of modality-specific ones to retain prior knowledge. Several methods fall under this umbrella [70, 17, 51], including prompt tuning, visual adapter, etc. Although effective, PEFT-based methods impose rigid constraints on the primary model weights, resulting in underfitting (i.e., a restricted upper bound on training performance) when handling the significant distribution drifts. In summary, existing fine-tuning paradigms fluctuate between excessive flexibility and over-restriction, both contributing to misfitting and a sub-optimal plasticity-stability trade-off between pre-trained knowledge and downstream adaptation (as Fig. 1).
In this work, we endeavor to mitigate the misfitting dilemma in cross-modal tracker adaptation by strategically regularizing the learning process. By probing parameter responses spanning from pre-training to fine-tuning phases, we identify that pronounced parameter sensitivities are the primary cause of degraded prior generalizability and transfer adaptability. To mitigate this, we propose a sensitivity-regularized fine-tuning framework that delicately calibrates the gradient updates to boost model transfer with precision. Specifically, we optimize multi-modality trackers from the following perspectives.
@@
    Formulating Prior Sensitivity. We commence by investigating the tangent space of pre-trained parameters, utilizing it as a critical indicator to assess and preserve prior generalizability for downstream fine-tuning. Furthermore, we estimate this sensitivity via an eigen-decomposition approximation.
@@
    Modeling Transfer Sensitivity. To further elucidate the adaptation challenges, we explore how sparse gradients exacerbate tuning instability. Accordingly, we formulate transfer sensitivity using off-the-shelf gradient matrices, aiming to facilitate effective gradient rebalancing.
@@
    Sensitivity Regularized Tuning. By harnessing the identified parameter sensitivities, we suggest an adaptive tuning scheme that safeguards essential pre-trained knowledge and fosters coherent multi-modality representations through finely modulated, sensitivity-driven updates. This mechanism facilitates seamless transfer across various multi-modality tracking tasks, continuously enhancing the model during the training phase.
Our method strategically guides the cross-domain fine-tuning process to optimize downstream tracking tasks. Extensive experimental results showcase our method achieves new state-of-the-art results across three multi-modality tracking tasks (RGB-Event, RGB-Depth, RGB-Thermal) and seven benchmarks, spanning diverse pre-trained tracking models. Comprehensive ablation studies demonstrate the effectiveness of the sensitivity-aware regularization fine-tuning concept.
@@
  • we revisit the misfitting issue of multi-modality tracking for adapting foundation models and propose a novel regularized tuning framework to indicate better transferability, which is orthogonal to the existing FFT and PEFT methods;
  • we formulate the parameter sensitivity with respect to pre-trained knowledge and transfer stability, and introduce a sensitivity-aware update strategy to refine the learning process, thereby facilitating the generalization and adaptability of trackers;
  • we conduct comprehensive experiments covering three multi-modality tracking tasks and seven benchmarks using diverse pre-trained trackers, and consistently push cross-modal tracking accuracy to new levels.
The remainder of the paper is organized as follows. Section II reviews existing literature on multi-modal trackers and cross-modal transfer learning methods. In Section III, we rigorously formulate prior and transfer sensitivities from both pre-training and fine-tuning perspectives, followed by the implementation of sensitivity-regularized tuning. Section IV presents extensive experiments to demonstrate the effectiveness of our method, along with comprehensive ablation studies to analyze the impact of different components and designs. Finally, Section V concludes the paper.
@@ II Related Work
II-A Multi-Modal Object Tracking
Object tracking, a cornerstone task in computer vision, involves predicting the position and scale of an object in future frames based on its initial appearance [48, 2, 28, 68]. However, the unreliability of RGB-only paradigms to adverse conditions has catalyzed research into multi-modal tracking, where auxiliary signals (e.g., event, depth, thermal data) are integrated to remedy the intrinsic deficiencies of visible imagery [66]. For instance, event cameras provide robust dynamic scene information under extreme motion or lighting. Leveraging this, [72, 11, 46, 63, 43, 44] combined RGB and event streams to predict objects in high-speed and low-dynamic scenarios. Meanwhile, depth cues offer geometric priors that are particularly useful for handling occlusion and background clutter. [32, 36, 56, 31] exploited this by incorporating depth maps into a tracking framework to improve robustness in crowded scenes. Similarly, thermal infrared imaging captures thermodynamic signatures independent of visible light, proves effective in low-illumination. [62, 19, 27, 52] demonstrated that fusing thermal and RGB data can yield more reliable appearance representations.
In summary, contemporary methods emphasized effective multi-modal interaction and fusion [42, 45, 59, 65]. With the emergence of large-scale RGB datasets and universal backbones (e.g., vision transformer), pre-trained trackers [58, 49, 9] have demonstrated remarkable generalization across diverse scenarios. These advances have shifted the paradigm toward transferring knowledge from pre-trained models in designing high-performance trackers. Consequently, multi-modal tracking is increasingly driven by reusing RGB-induced semantic priors, particularly when annotated multi-modal data are scarce or noisy. In this work, we target this imperative by optimizing the adaptation of pre-trained trackers for efficacious cross-modality transfer.
@@ II-B Cross-modal Transfer Learning
To adapt pre-trained models for multi-modal tracking, two main transfer learning strategies have recently emerged. Some works adhere to the full fine-tuning (FFT) paradigm [42, 45, 73, 41], where pre-trained models serve as weight initializations and are entirely re-trained on the target tasks. These methods require a shared or compact cross-modal feature space to inherit the generalization capability of the original model. Representatively, SUTrack [9] employed a single vision transformer with unified input representations for multiple modalities, avoiding task-specific customization. While effective, one primary dilemma may be innate to FFT: the contradiction between the paucity of large-scale downstream datasets and the huge appetites of cross-domain adaptation. This imbalance frequently leads to catastrophic forgetting of the pre-trained knowledge and severe overfitting to the limited target data. To alleviate this, profiting from the affluent experience of natural language processing and computer vision communities [21, 7, 18], the field has pivoted toward parameter-efficient fine-tuning (PEFT). The core principle of PEFT is to keep the majority of pre-trained weights frozen while tuning only a minimal number of additional parameters dedicated to the new task. By lightly fine-tuning, these methods aim to preserve the generalization while mitigating overfitting. For example, ProTrack [57] and ViPT [70] modulated RGB features by introducing trainable auxiliary tokens into attention layers. Similarly, methods like BAT [5] and SDSTrack [17] inserted lightweight adapter modules between attention layers for cross-modal shift compensation. More recently, unified trackers like UnTrack [51] and OneTracker [16] employed LoRA and prompt tuning to seamlessly integrate multiple modalities, enabling effective unification across diverse inputs. Despite their effectiveness, PEFT methods suffer from intrinsic limitations. First, the rigid constraints can lead to underfitting, limiting the model’s capacity to handle vast distribution drifts. Conversely, the randomly initialized add-on modules lack alignment with the pre-trained manifold, creating an optimization gap that predisposes the model to overfitting. Consequently, achieving robust and reliable fine-tuning for multi-modal tracking remains an elusive goal.
Remark. While recent works have explored sensitivity-aware mechanisms in efficient fine-tuning (SPT) and continual learning (CL), our SRFT introduces a fundamentally distinct paradigm in objectives and mechanisms. Specifically, methods like SPT [15] and CDRA-SPT [8] utilized sensitivity to prioritize structural adaptation, such as parameter masking or dynamic rank allocation for parameter efficiency. Conversely, MACL [47] employed sensitivity to prevent forgetting in sequential tasks by minimizing loss variance. Crucially, these approaches typically operate with an exclusive focus on the target or current data. Diverging from this, SRFT establishes a gradient-regularized paradigm. We formulate sensitivity to balance the preservation of pre-trained knowledge with sparse transfer gradient, via a hybrid sensitivity from pre-trained to multi-modal contexts. This distinction renders our method orthogonal to existing FFT, PEFT, and SPT frameworks.
Figure 2: Network architecture of the multi-modality trackers. All modules are initialized with the weights of a pre-trained RGB tracker. We only fine-tune backbones and fusion blocks with sensitivity-aware regularization terms.
@@ III Proposed Method
Learning generalized and coherent representations is crucial for adapting RGB-based models to multi-modal trackers. To unlock the full potential of pre-trained trackers, we revisit the core principles governing the cross-modal transfer process. In Section III-A, we first present the architectural design of our multi-modal trackers and highlight the key challenges associated with full fine-tuning. Next, Section III-B identifies and quantifies pronounced prior and transfer sensitivities that characterize pre-trained knowledge and cross-domain adaptation. Finally, in Section III-C, we incorporate these sensitivities into the optimization phase, enabling adaptive and dynamic modulation of parameter updates.
@@ III-A Preliminaries
Network Architecture of Multi-modality Trackers. Fig. 2 depicts the overall architecture of our multi-modal tracker, in which all modules are initialized with the weights of a pre-trained RGB tracker. For a multi-modal tracker, the RGB and auxiliary inputs are first fed to the embedding layer to generate the corresponding tokens. Then, symmetric transformer backbones (e.g., ViT or its variants) handle feature extraction and interaction. Without involving customized multi-modal fusion modules, we repurpose certain ViT blocks for multi-stage fusion of multi-modal tokens. Finally, the fused features are fed into the box head to estimate the object state. To retain modal-agnostic object association knowledge, the pre-trained box head is utilized and kept frozen.
Learning Process of Cross-Domain Transfer. We consider $f_{\theta}(\cdot):X\rightarrow Y$ as a multi-modal tracker with parameters $\theta\in\mathbb{R}^{|\theta|}$, with $|\theta|$ representing the total number of the parameters $\theta$. It takes task dataset $D=\{X,Y\}=\left\{\left(x_{i},y_{i}\right)\right\}_{i=1}^{M}$ to yield the optimal $\theta$, where $x$ represents a multi-modal input pair and $y$ is the target bounding box. In this paper, we focus on efficiently transferring the models pre-trained on a source domain $D_{0}$ (e.g., large-scale RGB data) to downstream domains $D_{t}$ (e.g., auxiliary modalities or degraded images), where typically $\left|D_{0}\right|\gg\left|D_{t}\right|$. In this setting, let $\theta_{0}$ be the pre-trained parameters (optimal on $D_{0}$), which serve as initialization. After fine-tuning on $D_{t}$, the parameters become $\theta_{t}$. Given a task-specific loss function $\mathcal{L}$, the vanilla transfer objective is:
$\theta_{t}=\underset{\theta}{\operatorname{\arg\min}}\mathcal{L}\left(\theta\mid D_{t}\right),\quad{s.t.}\quad\theta^{(0)}=\theta_{0},$ (1)
However, directly optimizing $\theta_{t}$ as per Eq. (1) leads to severe overfitting and unstable tuning, as evidenced by the pronounced train-test performance gap and oscillations in training dynamics observed in Fig. 1. This motivates a more principled transfer learning approach, as described next.
Figure 3: Loss-parameter manifold schematic across different fine-tuning paradigms. FFT updates all weights without constraints, risking severe forgetting of pre-trained knowledge (deteriorating pre-trained loss, L1). While PEFT restricts updates to additional parameters, this limits performance on new domains (stalled transferred loss, L2). In contrast, our SRFT performs optimization within an approximated pre-trained tangent space, indicating a better plasticity-stability trade-off.
III-B Modeling Sensitivity for Multi-Modality Tracking